A jar contains 99% fair coins, and 1% double-headed coins. you took a coin at random, then toss it 3 times and observed 3 heads. what is the probability that you picked a double-headed coin?

Verified Answer :

1 in 100, because you have 99% fair coins and just 1% double-headed one, then 99 + 1 is obviously 100, so you have a 1 in 100 chance

toporc

The probability of getting 3 heads in 3 tosses of a fair coin is:
\(0.5\times0.5\times0.5=0.125\)
The probability of selecting a fair coin and getting 3 heads in 3 tosses is:
\(0.99\times0.125=0.12375\)

The probability of selecting a double-headed coin and getting 3 heads in 3 tosses is:
\(0.01\times1=0.01\)
The probability of selecting a coin and getting 3 heads in 3 tosses is:
0.12375 + 0.01 = 0.13375
The probability that a double-headed coin was picked is:
0.01/0.13375 = 0.0748

The answer is 0.0748 or 7.48%